OpenStudy (anonymous):

Relativistic energy is sometimes derived in the following manner and I'm having trouble deciphering one mathematical step. So it goes like this: Given that \[F=\frac{ dp }{ dt }\] and reletavistic momentum is given by p=\frac{ mv }{ sqrt(1-\frac{v^2}{c^2}) } W=int_{x _{1}}^{x_{2}} F dx=int_{x _{1}}^{x_{2}} \frac{dp}{dt}dx which is fine then they say \frac{dp}{dt}=\frac{d}{dt}\frac{mv}{sqrt(1-(\frac{v^2}{c^2}))}=\frac{m(dv/dt)}{(1-(\frac{v^2}{c^2}))^\frac{3}{2}} its that last step that i dont understand. why does the denominator term now have a 3/2 index? and why is du/dt given in such an odd way?

OpenStudy (theeric):

Hi! I don't know if I'll be able to help or not, but I want to repost your question here with the equations in brackets. I don't see them appearing right now. "Relativistic energy is sometimes derived in the following manner and I'm having trouble deciphering one mathematical step. So it goes like this: Given that \[\large F=\frac{dp}{ dt}\] and reletavistic momentum is given by \[\large p=\frac{ mv }{\sqrt{1-\frac{v^2}{c^2}} }\] \[\large W=\int_{x _{1}}^{x_{2}} F\ dx=\int_{x _{1}}^{x_{2}} \frac{dp}{dt}\ dx\] which is fine then they say \[\large\frac{dp}{dt}=\frac{d}{dt}\frac{mv}{\sqrt{1-(\frac{v^2}{c^2})}}=\frac{m(dv/dt)}{(1-(\frac{v^2}{c^2}))^\frac{3}{2}}\] its that last step that i dont understand. why does the denominator term now have a 3/2 index? and why is du/dt given in such an odd way?"

OpenStudy (anonymous):

Thanks didnt know why they weren't showing up right.

OpenStudy (theeric):

Yeah! You typed it in well! And I'm wondering the same thing as you. I'm not a math buff, but it looks like all substitution until that last step, then they multiply the equation by \(\gamma^2\) and turn \(v\) into \(\large\frac{dv}{dt}\), which is acceleration, I think. So I don't know how they got that, at the moment. I didn't use that work equation at all, maybe that has something to do with it?

OpenStudy (theeric):

Well, I guess what they did in that last part is take the derivative with respect to \(t\). I missed that. I'm not sure if I can check that out with my own math! When my class had these equations and math processes, some of us just watched in awe and confusion... Others were better!

OpenStudy (theeric):

Did they take the derivative? Wow, I need to polish my math skills...

OpenStudy (anonymous):

It is integration by substitution. They then take the new expression for dp/dt and substitute it into the into to evaluate it.

OpenStudy (theeric):

\[\large\frac{d}{dt}\frac{mv}{\left(1-\left(\frac{v^2}{c^2}\right)\right)^\frac{1}{2}}=\frac{m\left(dv/dt\right)}{\left(1-\left(\frac{v^2}{c^2}\right)\right)^\frac{3}{2}}\] Does this mean the \(v\) gets differentiated with respect to \(t\)? Or does that come from somewhere else? \[\large\frac{d}{dt}\frac{m\boldsymbol v}{\left(1-\left(\frac{v^2}{c^2}\right)\right)^\frac{1}{2}}=\frac{m\boldsymbol{\left(dv/dt\right)}}{\left(1-\left(\frac{v^2}{c^2}\right)\right)^\frac{3}{2}}\] I'm sorry I can't be of more help. Good luck!

OpenStudy (egenriether):

Hey, there was too much writing to put this in the post so I attached the answer to a handwritten page I scanned in. Let me know if you need any clarification.

OpenStudy (anonymous):

Thanks